OpenStudy (anonymous):

Which sequence can be represented by the expression 12n? 144, 132, 120, 108,… 1, 12, 24, 36,… 12, 24, 36, 48,… 6, 18, 30, 42,…

OpenStudy (lalaly):

the third

OpenStudy (lalaly):

assuming n= 1,2,3,4,5,... 12*1 = 12 12*2= 24 12*3 = 36 12*4 =48 and so on